OpenStudy (anonymous):

Change 1.87 radians to degree measure. Round to the nearest tenth.

OpenStudy (johnweldon1993):

To convert from radians to degrees...you take your radian measure...and multiply it by \[\large \frac{180^\circ}{\pi}\] so we have \[\large 1.87 \times \frac{180^\circ}{\pi}\]